# Conch: Competitive Debate Analysis via Visualizing Clash Points and Hierarchical Strategies

### v1

@@ 3 Data Abstraction and Processing
As shown in Table 1, we collected a dataset from three representative competitive debates: the top-tier Chinese debate competition - International Chinese Debate Invitational [^2] https://bilibili.com/video/BV1Ca4y1y72t/ (ICDI), the largest academic competition - National Speech and Debate Tournament [^3] https://youtu.be/ytA7V3wXux0/ (NSDT), and the most prestigious debating society - the Oxford Union [^4] https://youtu.be/ByrhYvbwJoA/ (TOU). This collection includes debates conducted in two languages (Chinese and English) and three formats, offering a cross-cultural perspective on debate structures and strategies. Notably, interactions are more frequent and direct in ICDI than other competitions.
@@ 4.1 Process View
The conch-like diagram (Conch: Competitive Debate Analysis via Visualizing Clash Points and Hierarchical StrategiesA1-a) is composed of a series of parallel Archimedean spirals segments. As shown in Fig. 2, each Archimedean spiral segment corresponds to a debate session, arranged clockwise in chronological order, with its arc length representing the session’s content length. Each white and black segment (Fig. 2F) within the Archimedean spiral represents a block, with its arc length indicating the content length and color encoding two sides. With all Archimedean spirals constructed in a vertically oriented polar coordinate system, their start points are positioned on the axis and serve as centers for a column of tangent circles (numbered 1 - 8 in Fig. 2) in the Session View. These circles connect the Process View and the Strategy View, which are designed to represent the session’s content length and two sides’ content proportions.
@@
To analyze the clash-points-based interactions among blocks (R3), a chordal diagram is integrated inside the Archimedean spirals (Conch: Competitive Debate Analysis via Visualizing Clash Points and Hierarchical StrategiesA1-a), where each filled chord illustrates an interaction over a clash point between two blocks by connecting their respective sessions. As shown in Fig. 2E, the start and end points of the chord on each session’s arc indicate the block’s position within that session. The color of the chord indicates whether the two connected parties are on the same side or different sides regarding the clash point. If the parties are from different sides for this clash point, the line will display two distinct colors. Each color occupies half of the line and corresponds to the color of one party’s side. Conversely, if both parties belong to the same side concerning this clash point, the entire line will be a single color, representing their side’s color. Users are allowed to filter chords over a certain clash point or block. The periphery of the chordal graph is an Archimedean spiral (Conch: Competitive Debate Analysis via Visualizing Clash Points and Hierarchical StrategiesA1-b and Conch: Competitive Debate Analysis via Visualizing Clash Points and Hierarchical StrategiesA1-c). As shown in Fig. 2F, the area near the pole (Conch: Competitive Debate Analysis via Visualizing Clash Points and Hierarchical StrategiesA1-b) is radially divided into multiple color-filled sections. Each section represents a clash points in this session, and its width indicates the percentage of its corresponding clash point. The remaining central spiral and its adjacent sector-like areas (Conch: Competitive Debate Analysis via Visualizing Clash Points and Hierarchical StrategiesA1-c) display detailed information about sessions (R2). Each sector-like area (Fig. 2F), representing a single session, is further divided into multiple blocks by the disagreements within the session. The area of each block is proportional to the number of blocks in the corresponding disagreement, while their order is not meaningful. Each block (Fig. 2F1) is then filled with a specific color to indicate the clash point under discussion. In each block, the theme of the disagreement is highlighted in a larger font, while the viewpoints of the opposing sides are juxtaposed in a smaller font. Additionally, the debate strategies used in the disagreement are depicted as icons in the innermost layer of the area. All content is organized into blocks and arranged along a curve. The font size in each block is maximized based on the available space and the block’s proportional size. While this approach makes efficient use of space, it can cause some reading difficulty [51]. To address this, we provide a popup that displays the full text content when the user hovers over a block, ensuring better readability. Notably, a debater might mention a clash point in a session, but it does not become part of an interaction with another debater. In these situations, the color for that clash point will appear in the middle ring (Conch: Competitive Debate Analysis via Visualizing Clash Points and Hierarchical StrategiesA1-b) to show the topic was present, but there will be no corresponding disagreement block in the outer diagram (Conch: Competitive Debate Analysis via Visualizing Clash Points and Hierarchical StrategiesA1-c). This reflects the reality of debates, where not every point raised leads to a direct conflict.
@@ 4.2 Strategy View
The augmented stacked bar diagram is the core component of the Strategy View (Conch: Competitive Debate Analysis via Visualizing Clash Points and Hierarchical StrategiesA2), elaborately designed to display detailed strategy usage. This diagram is positioned alongside the Session View (Conch: Competitive Debate Analysis via Visualizing Clash Points and Hierarchical StrategiesA0) with each of its rows directly corresponding to a circle in the Session View (Conch: Competitive Debate Analysis via Visualizing Clash Points and Hierarchical StrategiesA0). The height of each row matches the circle’s diameter, visually representing the length of that session’s content. The diagram’s columns represent different debate strategy types, and the width of each column is significant: it indicates the highest frequency (peak usage) of that particular strategy across all sessions (Fig. 3G). These columns are arranged from left to right, starting with the strategy that has the lowest peak usage (narrowest column) and progressing to the one with the highest (widest column). Within this structure, each small rectangle, signifies a single instance of a strategy being used. The color of the unit clearly indicates which debate side used the strategy, while its position in a column tells us which specific strategy it was; an icon at the top of each column further identifies the strategy type. To visualize co-occurrence, if a block involves multiple strategies, the units representing these strategies are connected by solid lines across their respective columns, as shown in Fig. 3H. The leftmost of these connected units then has a dashed line extending to a dedicated area on the diagram’s left. This area displays icons representing all strategies that co-occurred in that specific block, with their colors also indicating the debate side. Furthermore, if the same set of co-occurring strategies appears multiple times within a session, their icon-based representations on the left may overlap. Each instance in the main bar chart will still link to this representation. Consequently, the number of dashed lines pointing to a particular box or overlapping area of co-occurring strategy icons directly reflects how often that specific combination of strategies was used together in the session.
@@
Design alternatives and Advantages. Common visualization methods for this type of data, such as stacked bar charts, mosaic graphs, and heat maps, each present certain limitations when trying to clearly display all necessary information. Traditional stacked bar charts (Fig. 4 a) make it difficult to accurately compare the sizes of different segments because they lack a common baseline, which also obscures the true differences between these segments. While mosaic graphs (Fig. 4 b) are more effective for showing proportions and the lengths of individual components, they still pose challenges for comparing components across different groups and tend to downplay the actual values involved. Similarly, heat maps (Fig. 4 c), though useful for illustrating the frequency of strategy use through color sequences, do not account for variations in debate duration, potentially leading to incorrect interpretations of the debate’s overall structure. These limitations highlighted several key challenges for our design: accurately showing strategy usage across different sessions, enabling easy comparison of strategy use both within a single session and between multiple sessions, and clearly representing differences in content length from one session to another. To address these challenges, we developed an augmented stacked bar diagram. This new design builds upon the familiar concept of a stacked bar chart but introduces two significant enhancements: a variable-width for the bars and non-full-format filling. The primary advantage of our approach is its ability to integrate multiple dimensions of information — including the content length of each session, the specific strategies employed, and the distribution of these strategy usages across all sessions — into a single, comprehensive view.
@@ 5.1 Case 1: Analyzing Debate Evolution and Interaction
E1 is the captain of a Chinese debate competition team, who has five years of debate experience. She was asked to analyze a Chinese debate competition from ICDI by Conch. The debate topic was “Whether Employment Prospects Should Be the Primary Consideration When Choosing University Majors After the Gaokao [^8] National College Entrance Examination (NCEE)”. Firstly, she explored the Overview Part, as shown in Conch: Competitive Debate Analysis via Visualizing Clash Points and Hierarchical StrategiesA. From the colorful filled blocks on the border of the chordal graph (Fig. 5(a)), she noticed that “Value Prioritization” was the main clash point of the debate and was discussed in all sessions, as indicated by the continuous appearance and large proportion of the red blocks. In addition, she found “Future Predictability” and “Decision-Making Approach” were also discussed continuously, with orange and yellow blocks appearing in multiple session areas. However, “Path to Happiness” was mainly discussed in the first half of the debate, as green blocks mainly appeared in the earlier area. Moreover, blocks in other colors were found appearing scattered across sessions, and therefore considered the corresponding clash point as non-core disagreements.
@@
Explore the clash points and the strategies. E1 noticed that yellow blocks appeared only in sessions involving the affirmative side (S1 in Fig. 5(b)). Therefore, she started focusing on the corresponding clash point — “Decision-Making Approach”. Based on this observation, she concluded that the affirmative side had an advantage on this clash point, as the negative side rarely participated in related discussions. Then, she used the strategy view (S2 in Fig. 5(b)) to explore the strategies used in relevant sessions. She noticed that black blocks, representing the negative side, appeared in the column marked with a handshake icon (“Refutation through Agreement”), but never appeared in the column marked with an eye icon (“Refutation through Ignoring”). This observation further confirmed her judgment, as the negative’s strategy was mainly to agree with the affirmative side. Consequently, she further analyzed the disagreements and related interactions of this clash point. She chose the three disagreements of this clash point in turn and observed their development path in the chordal graph. As shown in Fig. 6(a), Fig. 6(b) and Fig. 6(c), E1 found the sequential clustering of yellow blocks, and these three disagreements were all initiated by the affirmative side, clearly evolved one after another over time. Specifically, the disagreement “Choice Method” appeared first, starting from session 1 and lasting for four sessions, which formed the earliest cluster of yellow blocks (Fig. 6(a)). Immediately afterward, the disagreement “Guidance Tool” became the main topic and was actively discussed until session 5, forming the second clear cluster (Fig. 6(b)). Finally, in the second half of the debate, the disagreement “Decision Steps” emerged as the primary focus, as it was heavily discussed during the final six sessions and formed the last cluster of yellow blocks (Fig. 6(c)).
@@
Delve into the interaction within a disagreement. By reviewing the three progressively developed disagreements and related viewpoints, E1 found the negative side hold the view that “people should first reflect on themselves to learn what they want to learn before choosing college majors”. Since “Decision Steps” was the last disagreement of this clash point, whoever had an advantage on this disagreement would determine who won the entire clash point. E1 then followed the path under “Decision Steps” (Fig. 6(c)) to examine the detailed content. She noticed that both the beginning and end of the path corresponded to white blocks, and a long black block existed in the middle of the path. She focused particularly on understanding these three corresponding blocks. Through careful examination, she found that the affirmative side presented this disagreement using detailed and solid data, while the negative side only briefly argued their position and explained their position by “ideal”. In the final block, the affirmative side pointed out the negative side’s flaw of inconsistent reasoning, criticizing them for shifting from “reflecting on oneself” to “idealism” above all. Ultimately, she concluded that the affirmative side won this clash point. Interestingly, she noted that the negative side’s initial suggestion of “reflecting on oneself” did not even serve as a valid rebuttal to the affirmative side. If a person is indeed materialistic, then “reflecting on oneself” would naturally lead them to choose their major based on employment opportunities, which clearly aligns with the affirmative side’s position. This suggests that the negative side itself has a flaw in its logic.
@@ 5.2 Case 2: Understanding Overall Debate tactics
E2 first noticed that black and white chords appeared separately in the chord diagram, with very few chords mixing black and white (Fig. 7U0). This indicated that the affirmative and negative sides mainly developed arguments internally within their own sides, rather than between the two sides, which means direct interactions and clashes between the two sides are limited. She found this phenomenon unusual. Therefore, she further looked into the clash point labeled “Interpretation of the motion”, represented by the red color sections that frequently appeared and occupied significant space (Fig. 7U1). She discovered that the two sides differed fundamentally in their understanding and definitions of the debate topic itself. This difference explained why there were relatively few direct confrontations between them.
@@
Given these observations, E2 shifted her attention to the strategy view (Fig. 7U2) to better understand the strategic framework of the entire debate. She noticed that black blocks appeared extensively across all columns, whereas white blocks were more concentrated in the columns labeled “Refutation through Agreement” and “Refutation through Reasoning”. Considering both sides’ differing definitions of the motion, she identified distinctly opposite patterns between the two teams: the affirmative side primarily used reasoning to indirectly strengthen their own definition, rarely directly refuting the negative side. In contrast, the negative side heavily relied on evidence to directly refute the affirmative side.
E2 further examined co-occurring strategies and discovered that for the black side (negative), “Refutation through Reasoning” frequently appeared together with “Refutation through Evidence” (Fig. 7U2-a, Fig. 7U2-c). On the other hand, for the white side (affirmative), “Refutation through Reasoning” was more often combined with “Refutation through Agreement” (Fig. 7U2-b). She concluded that these insights further highlighted the systematic differences between the affirmative side and negative side throughout the debate, not only on basic interpretations, but also on strategic framework.

### v2

@@ 3 Data Abstraction and Processing
As shown in Table˜1, we collected a dataset from three representative competitive debates: the top-tier Chinese debate competition - International Chinese Debate Invitational [^2] https://bilibili.com/video/BV1Ca4y1y72t/ (ICDI), the largest academic competition - National Speech and Debate Tournament [^3] https://youtu.be/ytA7V3wXux0/ (NSDT), and the most prestigious debating society - the Oxford Union [^4] https://youtu.be/ByrhYvbwJoA/ (TOU). This collection includes debates conducted in two languages (Chinese and English) and three formats, offering a cross-cultural perspective on debate structures and strategies. Notably, interactions are more frequent and direct in ICDI than other competitions.
@@ 4.1 Process View
The conch-like diagram (Conch: Competitive Debate Analysis via Visualizing Clash Points and Hierarchical StrategiesA1-a) is composed of a series of parallel Archimedean spirals segments. As shown in Fig.˜2, each Archimedean spiral segment corresponds to a debate session, arranged clockwise in chronological order, with its arc length representing the session’s content length. Each white and black segment (Fig.˜2F) within the Archimedean spiral represents a block, with its arc length indicating the content length and color encoding two sides. With all Archimedean spirals constructed in a vertically oriented polar coordinate system, their start points are positioned on the axis and serve as centers for a column of tangent circles (numbered 1 - 8 in Fig.˜2) in the Session View. These circles connect the Process View and the Strategy View, which are designed to represent the session’s content length and two sides’ content proportions.
@@
To analyze the clash-points-based interactions among blocks (R3), a chordal diagram is integrated inside the Archimedean spirals (Conch: Competitive Debate Analysis via Visualizing Clash Points and Hierarchical StrategiesA1-a), where each filled chord illustrates an interaction over a clash point between two blocks by connecting their respective sessions. As shown in Fig.˜2E, the start and end points of the chord on each session’s arc indicate the block’s position within that session. The color of the chord indicates whether the two connected parties are on the same side or different sides regarding the clash point. If the parties are from different sides for this clash point, the line will display two distinct colors. Each color occupies half of the line and corresponds to the color of one party’s side. Conversely, if both parties belong to the same side concerning this clash point, the entire line will be a single color, representing their side’s color. Users are allowed to filter chords over a certain clash point or block. The periphery of the chordal graph is an Archimedean spiral (Conch: Competitive Debate Analysis via Visualizing Clash Points and Hierarchical StrategiesA1-b and Conch: Competitive Debate Analysis via Visualizing Clash Points and Hierarchical StrategiesA1-c). As shown in Fig.˜2F, the area near the pole (Conch: Competitive Debate Analysis via Visualizing Clash Points and Hierarchical StrategiesA1-b) is radially divided into multiple color-filled sections. Each section represents a clash point in this session, and its width indicates the percentage of its corresponding clash point. The remaining central spiral and its adjacent sector-like areas (Conch: Competitive Debate Analysis via Visualizing Clash Points and Hierarchical StrategiesA1-c) display detailed information about sessions (R2). Each sector-like area (Fig.˜2F), representing a single session, is further divided into multiple blocks by the disagreements within the session. The area of each block is proportional to the number of blocks in the corresponding disagreement, while their order is not meaningful. Each block (Fig.˜2F1) is then filled with a specific color to indicate the clash point under discussion. In each block, the theme of the disagreement is highlighted in a larger font, while the viewpoints of the opposing sides are juxtaposed in a smaller font. Additionally, the debate strategies used in the disagreement are depicted as icons in the innermost layer of the area. All content is organized into blocks and arranged along a curve. The font size in each block is maximized based on the available space and the block’s proportional size. While this approach makes efficient use of space, it can cause some reading difficulty [51]. To address this, we provide a popup that displays the full text content when the user hovers over a block, ensuring better readability. Notably, a debater might mention a clash point in a session, but it does not become part of an interaction with another debater. In these situations, the color for that clash point will appear in the middle ring (Conch: Competitive Debate Analysis via Visualizing Clash Points and Hierarchical StrategiesA1-b) to show the topic was present, but there will be no corresponding disagreement block in the outer diagram (Conch: Competitive Debate Analysis via Visualizing Clash Points and Hierarchical StrategiesA1-c). This reflects the reality of debates, where not every point raised leads to a direct conflict.
@@ 4.2 Strategy View
The augmented stacked bar diagram is the core component of the Strategy View (Conch: Competitive Debate Analysis via Visualizing Clash Points and Hierarchical StrategiesA2), elaborately designed to display detailed strategy usage. This diagram is positioned alongside the Session View (Conch: Competitive Debate Analysis via Visualizing Clash Points and Hierarchical StrategiesA0) with each of its rows directly corresponding to a circle in the Session View (Conch: Competitive Debate Analysis via Visualizing Clash Points and Hierarchical StrategiesA0). The height of each row matches the circle’s diameter, visually representing the length of that session’s content. The diagram’s columns represent different debate strategy types, and the width of each column is significant: it indicates the highest frequency (peak usage) of that particular strategy across all sessions (Fig.˜3G). These columns are arranged from left to right, starting with the strategy that has the lowest peak usage (narrowest column) and progressing to the one with the highest (widest column). Within this structure, each small rectangle, signifies a single instance of a strategy being used. The color of the unit clearly indicates which debate side used the strategy, while its position in a column tells us which specific strategy it was; an icon at the top of each column further identifies the strategy type. To visualize co-occurrence, if a block involves multiple strategies, the units representing these strategies are connected by solid lines across their respective columns, as shown in Fig.˜3H. The leftmost of these connected units then has a dashed line extending to a dedicated area on the diagram’s left. This area displays icons representing all strategies that co-occurred in that specific block, with their colors also indicating the debate side. Furthermore, if the same set of co-occurring strategies appears multiple times within a session, their icon-based representations on the left may overlap. Each instance in the main bar chart will still link to this representation. Consequently, the number of dashed lines pointing to a particular box or overlapping area of co-occurring strategy icons directly reflects how often that specific combination of strategies was used together in the session.
@@
Design alternatives and Advantages. Common visualization methods for this type of data, such as stacked bar charts, mosaic graphs, and heat maps, each present certain limitations when trying to clearly display all necessary information. Traditional stacked bar charts (Fig.˜4 a) make it difficult to accurately compare the sizes of different segments because they lack a common baseline, which also obscures the true differences between these segments. While mosaic graphs (Fig.˜4 b) are more effective for showing proportions and the lengths of individual components, they still pose challenges for comparing components across different groups and tend to downplay the actual values involved. Similarly, heat maps (Fig.˜4 c), though useful for illustrating the frequency of strategy use through color sequences, do not account for variations in debate duration, potentially leading to incorrect interpretations of the debate’s overall structure. These limitations highlighted several key challenges for our design: accurately showing strategy usage across different sessions, enabling easy comparison of strategy use both within a single session and between multiple sessions, and clearly representing differences in content length from one session to another. To address these challenges, we developed an augmented stacked bar diagram. This new design builds upon the familiar concept of a stacked bar chart but introduces two significant enhancements: a variable-width for the bars and non-full-format filling. The primary advantage of our approach is its ability to integrate multiple dimensions of information — including the content length of each session, the specific strategies employed, and the distribution of these strategy usages across all sessions — into a single, comprehensive view.
@@ 5.1 Case 1: Analyzing Debate Evolution and Interaction
E1 is the captain of a Chinese debate competition team, who has five years of debate experience. She was asked to analyze a Chinese debate competition from ICDI by Conch. The debate topic was “Whether Employment Prospects Should Be the Primary Consideration When Choosing University Majors After the Gaokao [^8] National College Entrance Examination (NCEE)”. Firstly, she explored the Overview Part, as shown in Conch: Competitive Debate Analysis via Visualizing Clash Points and Hierarchical StrategiesA. From the colorful filled blocks on the border of the chordal graph (Fig.˜5(a)), she noticed that “Value Prioritization” was the main clash point of the debate and was discussed in all sessions, as indicated by the continuous appearance and large proportion of the red blocks. In addition, she found “Future Predictability” and “Decision-Making Approach” were also discussed continuously, with orange and yellow blocks appearing in multiple session areas. However, “Path to Happiness” was mainly discussed in the first half of the debate, as green blocks mainly appeared in the earlier area. Moreover, blocks in other colors were found appearing scattered across sessions, and therefore considered the corresponding clash point as non-core disagreements.
@@
Explore the clash points and the strategies. E1 noticed that yellow blocks appeared only in sessions involving the affirmative side (S1 in Fig.˜5(b)). Therefore, she started focusing on the corresponding clash point — “Decision-Making Approach”. Based on this observation, she concluded that the affirmative side had an advantage on this clash point, as the negative side rarely participated in related discussions. Then, she used the strategy view (S2 in Fig.˜5(b)) to explore the strategies used in relevant sessions. She noticed that black blocks, representing the negative side, appeared in the column marked with a handshake icon (“Refutation through Agreement”), but never appeared in the column marked with an eye icon (“Refutation through Ignoring”). This observation further confirmed her judgment, as the negative’s strategy was mainly to agree with the affirmative side. Consequently, she further analyzed the disagreements and related interactions of this clash point. She chose the three disagreements of this clash point in turn and observed their development path in the chordal graph. As shown in Fig.˜6(a), Fig.˜6(b) and Fig.˜6(c), E1 found the sequential clustering of yellow blocks, and these three disagreements were all initiated by the affirmative side, clearly evolved one after another over time. Specifically, the disagreement “Choice Method” appeared first, starting from session 1 and lasting for four sessions, which formed the earliest cluster of yellow blocks (Fig.˜6(a)). Immediately afterward, the disagreement “Guidance Tool” became the main topic and was actively discussed until session 5, forming the second clear cluster (Fig.˜6(b)). Finally, in the second half of the debate, the disagreement “Decision Steps” emerged as the primary focus, as it was heavily discussed during the final six sessions and formed the last cluster of yellow blocks (Fig.˜6(c)).
@@
Delve into the interaction within a disagreement. By reviewing the three progressively developed disagreements and related viewpoints, E1 found the negative side hold the view that “people should first reflect on themselves to learn what they want to learn before choosing college majors”. Since “Decision Steps” was the last disagreement of this clash point, whoever had an advantage on this disagreement would determine who won the entire clash point. E1 then followed the path under “Decision Steps” (Fig.˜6(c)) to examine the detailed content. She noticed that both the beginning and end of the path corresponded to white blocks, and a long black block existed in the middle of the path. She focused particularly on understanding these three corresponding blocks. Through careful examination, she found that the affirmative side presented this disagreement using detailed and solid data, while the negative side only briefly argued their position and explained their position by “ideal”. In the final block, the affirmative side pointed out the negative side’s flaw of inconsistent reasoning, criticizing them for shifting from “reflecting on oneself” to “idealism” above all. Ultimately, she concluded that the affirmative side won this clash point. Interestingly, she noted that the negative side’s initial suggestion of “reflecting on oneself” did not even serve as a valid rebuttal to the affirmative side. If a person is indeed materialistic, then “reflecting on oneself” would naturally lead them to choose their major based on employment opportunities, which clearly aligns with the affirmative side’s position. This suggests that the negative side itself has a flaw in its logic.
@@ 5.2 Case 2: Understanding Overall Debate tactics
E2 first noticed that black and white chords appeared separately in the chord diagram, with very few chords mixing black and white (Fig.˜7U0). This indicated that the affirmative and negative sides mainly developed arguments internally within their own sides, rather than between the two sides, which means direct interactions and clashes between the two sides are limited. She found this phenomenon unusual. Therefore, she further looked into the clash point labeled “Interpretation of the motion”, represented by the red color sections that frequently appeared and occupied significant space (Fig.˜7U1). She discovered that the two sides differed fundamentally in their understanding and definitions of the debate topic itself. This difference explained why there were relatively few direct confrontations between them.
@@
Given these observations, E2 shifted her attention to the strategy view (Fig.˜7U2) to better understand the strategic framework of the entire debate. She noticed that black blocks appeared extensively across all columns, whereas white blocks were more concentrated in the columns labeled “Refutation through Agreement” and “Refutation through Reasoning”. Considering both sides’ differing definitions of the motion, she identified distinctly opposite patterns between the two teams: the affirmative side primarily used reasoning to indirectly strengthen their own definition, rarely directly refuting the negative side. In contrast, the negative side heavily relied on evidence to directly refute the affirmative side.
E2 further examined co-occurring strategies and discovered that for the black side (negative), “Refutation through Reasoning” frequently appeared together with “Refutation through Evidence” (Fig.˜7U2-a, Fig.˜7U2-c). On the other hand, for the white side (affirmative), “Refutation through Reasoning” was more often combined with “Refutation through Agreement” (Fig.˜7U2-b). She concluded that these insights further highlighted the systematic differences between the affirmative side and negative side throughout the debate, not only on basic interpretations, but also on strategic framework.